OpenStudy (anonymous):

find the exact values of the sine, cosine, and tangent of the angle (-105°) by using a sum or difference formula.

terenzreignz (terenzreignz):

Well, for starters sin(-x) = -sin(x) cos(-x) = cos(x) tan(-x) = -tan(x)